# Singular Value Scaling: Efficient Generative Model Compression via Pruned Weights Refinement

### v2

@@ Introduction
Diff-Prune 5.49 0.931 0.671 0.742 0.772 0.888 160K
Diff-Prune (scaled) 5.14 0.923 0.668 0.740 0.768 0.889
Diff-Prune 50% 8.96M 1.5G 7.77 0.914 0.670 0.720 0.805 0.861 360K
Diff-Prune (scaled) 7.41 0.912 0.671 0.718 0.817 0.868
Diff-Prune 70% 5.12M 1.0G 9.87 0.906 0.665 0.702 0.818 0.836 600K
Diff-Prune (scaled) 9.38 0.907 0.671 0.699 0.834 0.844
@@
Diff-Prune 6.17 0.949 0.591 0.803 0.502 0.760 100K
Diff-Prune (scaled) 5.57 0.938 0.619 0.793 0.573 0.785
Diff-Prune 50% 19.7M 6.0G 5.32 0.926 0.618 0.773 0.594 0.793 200K
Diff-Prune (scaled) 4.66 0.921 0.634 0.766 0.626 0.812
Diff-Prune 70% 9.25M 4.2G 6.30 0.913 0.584 0.772 0.524 0.761 240K
Diff-Prune (scaled) 5.04 0.905 0.625 0.752 0.631 0.812
@@
Table 3 presents the results of DDPM compression. For DDPM compression, we compare our method with the best-performing baseline during fine-tuning. Diff-Prune achieves higher SSIM scores, indicating closer alignment with the pre-trained model. This is expected, as our method refines pruned weights to balance knowledge preservation with easier fine-tuning. Consequently, our method consistently achieves better FID scores across all pruning ratios, indicating more effective distribution matching. Additionally, the modest drop in SSIM indicates that our method still effectively retains contextual knowledge, even as it further relaxes the knowledge of the pre-trained model.

### v3

@@ Introduction
Diff-Prune 5.49 0.931 0.742 0.671 0.935 0.875 160K
Diff-Prune (scaled) 5.14 0.923 0.740 0.668 0.936 0.872
Diff-Prune 50% 8.96M 1.5G 7.77 0.914 0.720 0.670 0.840 0.815 360K
Diff-Prune (scaled) 7.41 0.912 0.718 0.671 0.836 0.816
Diff-Prune 70% 5.12M 1.0G 9.87 0.906 0.702 0.665 0.777 0.767 600K
Diff-Prune (scaled) 9.38 0.907 0.699 0.670 0.774 0.772
@@
Diff-Prune 6.17 0.949 0.803 0.591 1.043 0.881 100K
Diff-Prune (scaled) 5.57 0.938 0.793 0.619 1.011 0.889
Diff-Prune 50% 19.7M 6.0G 5.32 0.926 0.773 0.618 0.928 0.875 200K
Diff-Prune (scaled) 4.66 0.921 0.766 0.634 0.925 0.885
Diff-Prune 70% 9.25M 4.2G 6.30 0.913 0.772 0.584 0.928 0.857 240K
Diff-Prune (scaled) 5.04 0.905 0.752 0.625 0.873 0.864
@@
Table 3 presents the results of DDPM compression. [^5] During the rebuttal phase, we found a mismeasurement in Precision, Recall, Density, and Coverage in Table 3. After correcting it, the results are now more consistent, reaffirming that Singular Value Scaling preserves diversity well. 1q‘ For DDPM compression, we compare our method with the best-performing baseline during fine-tuning. Diff-Prune achieves higher SSIM scores, indicating closer alignment with the pre-trained model. This is expected, as our method refines pruned weights to balance knowledge preservation with easier fine-tuning. Consequently, our method consistently achieves better FID scores across all pruning ratios, indicating more effective distribution matching. Additionally, the modest drop in SSIM indicates that our method still effectively retains contextual knowledge, even as it further relaxes the knowledge of the pre-trained model.